# Hyperdimensional Computing for Node Classification and Link Prediction

### v2

@@ 4.3 Node Label Prediction using HDGL
With the procedure by which HDGL constructs the HD representation for each of the nodes in a graph, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$. Here, we follow the standard approach to predicting class labels of HD encoded objects [15]. Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes. During the inference phase, given the HD vectors for each of the classes, we predict the label of an unlabeled node $u$ as simply the label associated with the class hyper-vector that is, among all class HD class vectors is closest to the HD vector $\mathbf{z}_{u}$, the encoding of $u$. More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph. The class HD are constructed for each class as follows: $\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$, resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$. Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted. We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows: $\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}]$.
4.4 Link Prediction Using HDGL
Section 4.1 and 4.2 outline a technique for constructing node representations optimized for node classification. Here we show that the same node representation can be used to support the transductive link prediction task.
Kanerva [22, 23] describes a method for encoding entries of a dataset. Let $\mathbf{a}_{1}$, $\mathbf{a}_{2}$, $\mathbf{a}_{3}$ be some randomly sampled HD vectors representing data entry attributes (eg. weight, height and gender). Let $\mathbf{\alpha}$, $\mathbf{\beta}$ and $\mathbf{\gamma}$ be values for the corresponding attributes represented in HD format. Let $\mathbf{r}$ be a HD vector such that: $\mathbf{r}=(\mathbf{a}_{1}\otimes\mathbf{\alpha})\oplus(\mathbf{a}_{2}\otimes\mathbf{\beta})\oplus(\mathbf{a}_{3}\otimes\mathbf{\gamma}).$ From the properties of HD-computing [22], it is possible to retrieve values for a corresponding attribute using the Binding $\otimes$ operator. Specifically, $\mathbf{\alpha}\approx\mathbf{r}\otimes\mathbf{a}_{1};\quad\mathbf{\beta}\approx\mathbf{r}\otimes\mathbf{a}_{2};\quad\mathbf{\gamma}\approx\mathbf{r}\otimes\mathbf{a}_{3};$
We apply the same concept to link prediction. Let $\mathbf{e}^{+}$ be a hypervector storing information about all or some of the edges present in the graph, and let $\mathbf{e}^{-}$ be a hypervector representing information about some of the non-edges in the graph. Formally: $\mathbf{e}^{+}=\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v});\quad\mathbf{e}^{-}=\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v}).$ Let $\mathbf{Z}$ have $N$ rows, storing latent representations of all the nodes in the graph. That is, a row in $\mathbf{Z}$ is represented by $\mathbf{z}_{u}$, where $u\in\mathcal{V}$. Now that we have encoded edges and non-edges in the graph, the goal is to find new or missing links in the graph. This can also be thought of as reconstructing/generating the adjacency matrix.
@@ 4.4 Link Prediction Using HDGL
To find possible new links we Bind $\otimes$ the edge information hypervector $\mathbf{e}^{+}$ and the latent representation of all the nodes in the graph. This results in a new matrix $\mathbf{Z}^{+}$ where each row is $\mathbf{z}_{u}\otimes\mathbf{e}^{+}$ where $u\in\mathcal{V}$. $\mathbf{Z}^{+}$ represents the possible node representations of the target nodes in a link, with the corresponding source node representation being $\mathbf{Z}$. We follow the same procedure for the non-edge hypervector, resulting in $\mathbf{Z}^{-}$. Precisely:
$\mathbf{Z}^{+}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{+}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{+}\end{bmatrix}$ $\mathbf{Z}^{-}$ $=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{-}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{-}\end{bmatrix}$
@@
Now, we find pairwise normalized distances between rows of $\mathbf{Z}^{+}$ and rows of $\mathbf{Z}$. This results in $\mathbf{D}^{+}=dist(\mathbf{Z}^{+},\mathbf{Z})\in[0,1]^{N\times N}$ , a matrix representing distances. If entry $(i,j)$ is close to zero, it indicates that the distance between row $i$ of $\mathbf{Z}^{+}$ is close to row $j$ in $\mathbf{Z}$, suggesting that $(i,j)$ might be a possible link. We perform the same process for $\mathbf{Z}^{-}$, resulting in the $\mathbf{D}^{-}=dist(\mathbf{Z}^{-},\mathbf{Z})\in[0,1]^{N\times N}$ matrix. However, an entry close to zero in $\mathbf{D}^{-}$ indicates a high possibility of not having an edge between the corresponding nodes.
We now construct a prediction of the adjacency matrix represented by $\hat{\mathbf{A}}$, where entries $\hat{A}_{ij}\in[0,1]$. To construct the final reconstructed or predicted adjacency matrix of probabilities (based on distances), which factors in both the distances of having, or not having an edge using $\mathbf{Z}^{+}$ and $\mathbf{Z}^{-}$, we compute:
@@ 5 Experiments
We proceed to report results of extensive experiments comparing HDGL with a set of state-of-the-art node classification methods and link prediction methods, including several GNN models, in terms of both predictive performance and runtime.
5.1 Data sets
Since HDGL is designed in such a way that nodes with similar features and neighborhoods will have similar representations, we exclusively concentrate on homophilous graphs in this study, deferring the exploration of more complex cases involving heterophily graphs to future research.
@@ 5.1 Data sets
For our experiments, we utilize commonly used node labeling homophilous graph benchmarks: the CORA graph dataset [61], sourced from McCallum et al. [32]; the CiteSeer graph dataset [61], sourced from Giles et al. [16]; and the PubMed dataset [36]. Additionally, we include three more datasets from the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al. [45], and the BlogCatalog Graph dataset from Huang et al. [20]. Additionally, we introduce a new version of the DBLP dataset, built upon the works of Pan et al. [39]. We remove nodes with no neighbors in the graph obtained from Pan et al. [39]. The original node features from Pan et al. [39] consist of manuscript titles. We utilize BERT [11] to obtain embeddings for the titles and employ them as node features. The data sets and their statistics are given in Table 1.
@@
We adhere to the Train/Validation/Test splits described in Yang et al. [61], Kipf and Welling [26], and Veličković et al. [53] for the CORA, CiteSeer, and PubMed datasets, respectively. For the Coauthor CS, Coauthor Physics, BlogCatalog, and DBLP datasets, we employ the data splitting strategy outlined in Shchur et al. [45].
We test link prediction models on CORA, CiteSeer, and Pubmed datasets. These models are trained on incomplete data, retaining node features but omitting some edges. Validation and test sets include excluded edges and randomly selected unconnected node pairs. Evaluation assesses the models’ ability to classify both edges and non-edges. We follow splitting strategy from Kipf and Welling [25] work i.e Validation and test sets each contain 5% and 10% of citation links.
5.2 Models
@@
For semi-supervised node classifcation task, we compare the performance of HDGL with that of several strong baselines Logistic Regression (LogReg), DeepWalk [42] and Label Propagation [64]. We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) [26], Graph Attention Network (GAT) [53], and lastly, a fast GNN model, namely, Simplified Graph Convolutional Network (SGC) [57] which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input. We also assess our model against RelHD [24], an existing HD algorithm used for node classification.
@@
For the transductive link prediction task, we evaluate our model against standard link prediction baselines:- DeepWalk [42] and Spectral clustering (SC) [48]. We also compare our model against graph autoencoder models:- Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) from Kipf and Welling [25].
5.3 HDGL Model Specifications
@@
Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 3.2), we opt to randomly sample an odd number, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL for datasets except BlogCatalog. For BlogCatalog we sample 21 of the 1-hop neighbors and 51 of the 2-hop neighbors since Blogcatalog graph is denser than the other datasets.
For node classification in CORA, CiteSeer, and BlogCatalog datasets, we map node features to $50,000$-dimensional HD vectors through random projections. Similarly, for the PubMed, Coauthor CS, Coauthor Physics, and DBLP datasets, we map node features to $20,000$-dimensional HD vectors using the same method. While these choices may seem arbitrary, it’s crucial to note that selecting dimensions between $20,000$ and $50,000$ typically suffices to maintain the desired properties of HD-computing. This assertion finds extensive support in the research from Kanerva [22, 23].
Table 2: Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of RelHD and HDGL Across Datasets (* denotes datasets with $\{0,1\}$ features. In these datasets, RelHD performs comparably to HDGL)
@@
5.4 Training and Evaluation
Recall that iterative machine learning methods, including GNN typically make use of a validation data set that is distinct from the test data set for deciding when to terminate training, or for tuning the hyper-parameters. Hence, effectively, they make use of both training and validation data sets for building the predictive model and test data for evaluating the trained model’s performance. In contrast, learning methods such as HDGL which invole no iterative training do not need validation data. Hence, to ensure fair comparison between HDGL and other methods, we evaluate their performance on the training data, while letting the individual methods use training and validation data as appropriate for the algorithm in question. Thus, HDGL uses training and validation sets together for learning, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning.
@@ 5.4 Training and Evaluation
In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data. For tuning the hyperparameters, we used grid search over a set of 12 configurations. In the case of GCN and GAT, the following hyperparameters were optimized the number of hidden units in the first layer, learning rate, and weight decay. In the case of GCN,networks with 32 and 64 hidden units were explored whereas in the case of GAT, 4 and 8 attention heads, each with 8 hidden units were explored. Both GCN and GAT were explored with weight decays in [1e-3, 1e-4, 5e-4]. In the case of SGC which does not use any hidden layers, only the learning rate and weight decay parameters were tuned in [1e-3, 5e-4, 1e-4, 5e-5, 1e-5, 5e-6]. GCN, GAT, and SGC were tuned for learning rates set to 1e-2 and 1e-3. The early stopping criterion optimized over choices of 25 and 50 epochs, respectively, for all three GNN models.
@@
As LogReg, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search. For Label Propagation, we use 20 iterations of propagation (except for blogcatalog dataset where we use 10) and set the alpha factor  [^1] Alpha factor in Label Propagation determines the balance between existing and propagated labels. Higher alpha prioritizes existing labels; lower alpha favors propagated ones. to 0.5. For DeepWalk, we utilize the DGL library, setting the embedding dimension to 32 and the number of walks to 20, with a batch size of 64. We sample such walks for all nodes in the graph and train DeepWalk model for 10 epochs and then employ logistic regression for classification, incorporating early stopping based on the validation set.
@@
Following Kang et al. [24], the RelHD algorithm was employed with 10,000 dimensional random basis HD vectors, as the performance of RelHD plateaus around this dimensionality.
@@
We report measured performance on the test data, averaged across 10 random selections of pertinent parameters. These parameters include the random hyperplanes employed for constructing high-dimensional representations used by HDGL, random basis vector initializations for RelHD, weight initializations of the deep learning models, and data splits for datasets such as Coauthor Physics, Coauthor CS, BlogCatalog, and DBLP. For Cora, Citeseer, and Pubmed, we use predefined splits given in Yang et al. [61].
@@
For the link prediction task, we rely on the results directly obtained from the paper for GVAE from Kipf and Welling [25].
5.5 Implementation Details
All experiments were conducted on Google Colab Pro with High-memory CPU. To ensure fair comparison, although not ideal for HDGL, we used the Python PyTorch framework [40], complemented by the Deep Graph Library (DGL) [54] for importing datasets and implementing deep learning models. None of our experiments used parallel execution or GPUs. We have also developed a PyTorch version of RelHD, as the codebase was not publicly available and experiments detailed in Kang et al. [24] were reliant on hardware architecture implementations. We undertake this effort to compare with our model and use RelHD as a baseline in this study.
5.6 Results
Node Classification Performance. The results of our comparison of HDGL with the other methods are shown in Table 2. We observe that HDGL without the need for computationally expensive iterative training or hyperparameter optimization achieves node prediction accuracy that is competitive with those of state-of-the-art GNN models (GCN, GAT, SGC) across most of the benchmark data sets, in each case reaching performance that is within 1-2% of the best performing model. HDGL also consistently outperforming basic deep learning baseline methods (LogReg, DeepWalk, LabelProp). The notable advantage of HDGL over RelHD lies in its ability to accommodate features beyond the binary $\{0,1\}$ range, as evident from the results presented in Table 2, where HDGL significantly outperforms. This is particularly notable with datasets lacking $\{0,1\}$ features. However, when dealing with datasets featuring $\{0,1\}$ attributes like CORA and Citeseer, the performance of HDGL and RelHD appears nearly identical.
To highlight the advantage of HDGL more prominently, we conducted experiments for node classification on the CORA and CiteSeer datasets where we reduced the node features to 100 dimensions using PCA [41], thereby trying to enforce continuous node features. As depicted in Figure 2, the impact is evident through a noticeable drop in performance for RelHD. Conversely, for other methods including HDGL, the performance remains relatively unchanged.
Figure 2: Impact of performing PCA with a dimensionality of 100 on features for node classification accuracy.
@@ 5.6 Results
Figure 3: Comparison of Learning Times (in seconds) i.e Across Datasets for HDGL and RelHD with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations. Runtime for HDGL and RelHD exclusively includes only time Elapsed for learning label representations using train/validation data.
Table 3: Link prediction results averaged over 10 Randomized Instantiations of HDGL and datasplits. Results for other models are from Kipf and Welling [25] paper.
Methods Cora Citeseer PubMed
AUC AP AUC AP AUC AP
DeepWalk 0.831 0.850 0.805 0.836 0.844 0.841
SC 0.846 0.885 0.791 0.826 0.849 0.888
GAE 0.910 0.920 0.895 0.899 0.964 0.965
VGAE 0.914 0.926 0.908 0.920 0.944 0.947
HDGL 0.849 0.880 0.768 0.842 0.853 0.884

### v3

@@ 4.3. Node Label Prediction using HDGL
With the procedure for constructing a HD representation of nodes and their local topologies in place, we turn to the task of using the resulting HD node representation to predict the node labels using $\mathbf{z}_{v}$. Here, we follow the standard approach to predicting class labels of HD encoded objects (Ge and Parhi, 2020). Specifically, during the learning phase, we Bundle the collection of HD encodings of nodes belonging to each class yielding a HD Class Hyper-Vector for each of the classes. During the inference phase, given the HD vectors for each of the classes, we predict the label of an unlabeled node $u$ as simply the label associated with the class hyper-vector that is, among all class HD class vectors is closest to the HD vector $\mathbf{z}_{u}$, the encoding of $u$. More precisely let $\forall\ell\in\{1,\cdots,L\}$, $\mathcal{T}_{\ell}\subset\mathcal{V}$ be the subset of nodes that make up the training set with label $\ell$ and $\mathcal{V}_{\text{test}}\subset\mathcal{V}$ be the subset of unlabeled nodes in the graph. The class HD vectors are constructed for each class as follows: $\mathbf{c}_{\ell}=\mathop{\oplus}_{i\in\mathcal{T}_{\ell}}\mathbf{z}_{i}\quad;\;\forall\ell\in\{1,\cdots,L\}$, resulting in a set of class HD vectors $\{\mathbf{c}_{1},\cdots,\mathbf{c}_{L}\}$. Let $u\in\mathcal{V}_{\text{test}}$ be a node for which the class label is to be predicted. We predict the class label for $u$ using its HD representation and the HD representations of each of the classes as follows: $\hat{y}_{u}=\mathop{\operatorname{argmin}}_{\ell\in\{1,\cdots,L\}}d_{H}[\mathbf{z}_{u},\mathbf{c}_{\ell}]$.
4.4. Link Prediction Using HDGL
Our approach takes advantage of a technique introduced by Kanerva (1988, 2009)’s for representing and retrieving data using HD representation. Lets consider a graph with 7 nodes and there are 3 edges in the graph betweem node $(1,2)$, $(3,6)$ and $(5,7)$. By binding latent representation between pairs of nodes which have edges between them and bundling the resulting vectors, we obtain a sum-vector that encapsulates the edge information of the graph: $\mathbf{e}^{+}=(\mathbf{z}_{1}\otimes\mathbf{z}_{2})\oplus(\mathbf{z}_{3}\otimes\mathbf{z}_{6})\oplus(\mathbf{z}_{5}\otimes\mathbf{z}_{7})$.
Edge information can be recovered solely using hypervectors $\mathbf{e}^{+}$ and $\mathbf{z}_{i};\forall i\in\mathcal{V}$. For eg. $\mathbf{z}_{2}\approx\mathbf{e}^{+}\otimes\mathbf{z}_{1}$ and $\mathbf{z}_{7}\approx\mathbf{e}^{+}\otimes\mathbf{z}_{5}$.
For the transductive link prediction problem, we leverage on the assumption that the underlying semantics governing existing links, which are based on latent representations of nodes in an edge, will also apply to new edges. For instance, if nodes $(1,2)$ are connected and we are predicting edges involving node $4$ and if $\mathbf{z}_{1}\approx\mathbf{z}_{4}$, then we would expect an edge $(4,2)$ to be predicted. Note that it’s also crucial to accurately predict non-edges to minimize false positives. Therefore, we also make use of hypervector $\mathbf{e}^{-}$ to represent information about edges that are absent in the graph.
Following this overview, we can now delve into a detailed explanation of our model. Formally: $\mathbf{e}^{+}=\mathop{\oplus}_{(u,v)\in\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v});\quad\mathbf{e}^{-}=\mathop{\oplus}_{(u,v)\notin\mathcal{E}}(\mathbf{z}_{u}\otimes\mathbf{z}_{v}).$ Given this encoding of edges that are present and the edges that are absent in the graph, we can proceed to predict links as follows:
@@ 4.4. Link Prediction Using HDGL
To find putative new links we Bind $\otimes$ the edge information hypervector $\mathbf{e}^{+}$ and the latent representation of all the nodes in the graph. This results in a new matrix $\mathbf{Z}^{+}$ where the $i^{th}$ row is $\mathbf{z}_{i}\otimes\mathbf{e}^{+}$; representing the possible node representation of the target node in a link, with the corresponding source node representation being $\mathbf{z}_{i}$. We follow the same procedure for the hypervector encoding information about the edges that are absent, resulting in $\mathbf{Z}^{-}$. Precisely:
(3) $\begin{array}[]{cc}\mathbf{Z}^{+}=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{+}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{+}\end{bmatrix}&\mathbf{Z}^{-}=\begin{bmatrix}\mathbf{z}_{1}\otimes\mathbf{e}^{-}\\ \vdots\\ \mathbf{z}_{N}\otimes\mathbf{e}^{-}\end{bmatrix}\end{array}$
@@
Now, we find pairwise normalized distances between rows of $\mathbf{Z}^{+}$ and rows of $\mathbf{Z}$. This results in $\mathbf{D}^{+}=dist(\mathbf{Z}^{+},\mathbf{Z})\in[0,1]^{N\times N}$ , a matrix representing distances. If entry $(i,j)$ is close to zero, it indicates that the ”distance” between row $i$ of $\mathbf{Z}^{+}$ is close to row $j$ in $\mathbf{Z}$, suggesting the possibility of a link $i$ and $j$. Similarly, we find pairwise normalized distances between rows of $\mathbf{Z}^{-}$ and rows of $\mathbf{Z}$ yielding $\mathbf{D}^{-}=dist(\mathbf{Z}^{-},\mathbf{Z})$. An entry close to zero in $\mathbf{D}^{-}$ suggests the absence of an edge between the corresponding nodes.
We now generate a prediction for the adjacency matrix, denoted by $\hat{\mathbf{A}}$, where $\hat{A}_{ij}\in[0,1]$. We predict values that fall within this interval because the evaluation metrics used are AUC-ROC and Average Precision. For real-world applications, an appropriate threshold can be selected depending on the specific use case to effectively interpret these predictions. To obtain the final predicted adjacency matrix of probabilities, which reflects the likelihood of an edge being present based on $\mathbf{D}^{+}$ and $\mathbf{D}^{-}$, we perform the following:
@@ 5. Experiments
We proceed to report results of extensive experiments comparing HDGL with a set of state-of-the-art node classification methods and link prediction methods, including several GNN models, in terms of both predictive performance and runtime. The code for HDGL can be found at: https://github.com/Abhishek-Dalvi410/HDGL.
5.1. Data sets
Since HDGL is designed to ensure that nodes with similar features and neighborhoods have comparable representations—often resulting in identical node labels—we focus exclusively on graphs with these characteristics as benchmark datasets for this study.
@@ 5.1. Data sets
For our experiments, we utilize commonly used node labeling graph benchmarks: the CORA graph dataset (Yang et al., 2016), sourced from McCallum et al. (2000); the CiteSeer graph dataset (Yang et al., 2016), sourced from Giles et al. (1998); and the PubMed dataset (Namata et al., 2012). Additionally, we include three more datasets from the Microsoft Academic Graph: Coauthor CS and Coauthor Physics from Shchur et al. (2018), and the BlogCatalog Graph dataset from Huang et al. (2017). Additionally, we introduce a new version of the DBLP dataset, built upon the works of Pan et al. (2016). We remove nodes with no neighbors in the graph obtained from Pan et al. (2016). The original node features from Pan et al. (2016) consist of manuscript titles. We utilize BERT (Devlin et al., 2019) to obtain embeddings for the titles and employ them as node features. The data sets and their statistics are given in Table 1.
@@
We adhere to the Train/Validation/Test splits described in Yang et al. (2016), Kipf and Welling (2017), and Veličković et al. (2018) for the CORA, CiteSeer, and PubMed datasets, respectively. For the Coauthor CS, Coauthor Physics, BlogCatalog, and DBLP datasets, we employ the data splitting strategy outlined in Shchur et al. (2018).
We test transductive link prediction models on CORA, CiteSeer, and Pubmed datasets and with train/val/test splits from Kipf and Welling (2016) work i.e validation and test sets each contain 5% and 10% of links.
Table 1. Dataset Summary.
Dataset Nodes Edges Features Feature Type
CORA 2708 5429 1433 $\{0,1\}$
CiteSeer 3327 4732 3703 $\{0,1\}$
PubMed 19717 44338 500 $(0,1)$
BlogCatalog 5196 171743 8189 $\mathbb{Z}^{+}$
Coauthor CS 18333 81894 6805 $\mathbb{Z}^{+}$
Coauthor Physics 34493 247962 8415 $\mathbb{Z}^{+}$
DBLP 17725 105781 768 $\mathbb{R}$
5.2. Models
@@
For semi-supervised node classifcation task, we compare the performance of HDGL with that of several strong baselines Logistic Regression (LogReg), DeepWalk (Perozzi et al., 2014) and Label Propagation (Zhu and Ghahramani, 2002). We also directly compare HDGL with state-of-the-art models: Graph Convolutional Network (GCN) (Kipf and Welling, 2017), Graph Attention Network (GAT) (Veličković et al., 2018), and lastly, a fast GNN model, namely, Simplified Graph Convolutional Network (SGC) (Wu et al., 2019) which is essentially a logistic regression model trained using features extracted from $k$-hop neighborhoods of nodes as input. We also assess our model against RelHD (Kang et al., 2022), an existing HD algorithm used for node classification
@@
For the transductive link prediction task, we evaluate our model against standard link prediction baselines:- DeepWalk (Perozzi et al., 2014) and Spectral clustering (SC) (Tang and Liu, 2011). We also compare our model against graph autoencoder models:- Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) from Kipf and Welling (2016).
5.3. HDGL Model Specifications
@@
Because Bundling of an even number of HD vectors can result in ties that would need to be broken using tie-breaking policy (See Section 3.2), we opt to randomly sample odd number of neighbors, specifically, 11 of the 1-hop neighbors and 21 of the 2-hop neighbors during the construction of HD node representations by HDGL for datasets except BlogCatalog. For BlogCatalog we sample 21 of the 1-hop neighbors and 51 of the 2-hop neighbors since Blogcatalog graph is denser than the other datasets.
For node classification in CORA, CiteSeer, and BlogCatalog datasets, we map node features to $50,000$-dimensional HD vectors through random projections. Similarly, for the PubMed, Coauthor CS, Coauthor Physics, and DBLP datasets, we map node features to $20,000$-dimensional HD vectors using the same method. While these choices may seem arbitrary, it’s crucial to note that selecting dimensions between $20,000$ and $50,000$ typically suffices to maintain the desired properties of HD-computing. This assertion finds extensive support in the research from Kanerva (1988, 2009).
Table 2. Average Test Set Accuracy and Standard Deviation (in Percentage) Over 10 Random Weight Initializations for Deep Learning Models and 10 Randomized Instantiations of RelHD and HDGL Across Datasets (* denotes datasets with $\{0,1\}$ features. In these datasets, RelHD performs comparably to HDGL)
@@
5.4. Training and Evaluation
Iterative machine learning methods like GNN use separate validation data to stop training and tune hyperparameters, alongside training data. In contrast, methods such as HDGL which involve no iterative training do not need validation data. Hence, to ensure fair comparison, HDGL uses training and validation sets together for learning, whereas GNN use training data for training the model, validation data for early stopping and hyperparameter tuning.
@@ 5.4. Training and Evaluation
In the case of GNN methods that rely on iterative training, we choose the models with hyperparameters tuned to obtain optimal performance on validation data. We used grid search over 12 different configurations for tuning these hyperparameters. For GCN and GAT, we optimized the following hyperparameters: the number of hidden units in the first layer, learning rate, and weight decay. Specifically, GCN models were tested with 32 and 64 hidden units, while GAT models were tested with 4 and 8 attention heads, each having 8 hidden units. Both GCN and GAT models were evaluated with weight decays of 1e-3, 1e-4, and 5e-4. SGC, which does not have hidden layers, had its hyperparameters tuned for learning rate and weight decay. The weight decay tuning of SGC was over the following values 1e-3, 5e-4, 1e-4, 5e-5, 1e-5, and 5e-6.
GCN, GAT, and SGC models were tuned with learning rates of 1e-2 and 1e-3, using early stopping criteria set to 25 epochs for the 1e-2 learning rate and 50 epochs for the 1e-3 learning rate.
@@
As LogReg, DeepWalk, and Label Prop represent well-established baselines, we do not conduct a hyperparameter search. For Label Propagation, we use 20 iterations of propagation (except for blogcatalog dataset where we use 10) and set the alpha factor  [^1] Alpha factor in Label Propa determines the balance between existing and propagated labels. Higher alpha prioritizes existing labels; lower alpha favors propagated ones. to 0.5. For DeepWalk, we utilize the DGL library, setting the embedding dimension to 32 and the number of walks to 20, with a batch size of 64. We sample such walks for all nodes in the graph and train DeepWalk model for 10 epochs and then employ logistic regression for classification, incorporating early stopping based on the validation set.
@@
Following Kang et al. (2022), the RelHD algorithm was employed with 10,000 dimensional random basis HD vectors, as the performance of RelHD plateaus around this dimensionality.
@@
We report measured performance on the test data, averaged across 10 random selections of pertinent parameters. These parameters include the random hyperplanes employed for constructing high-dimensional representations used by HDGL, random basis vector initializations for RelHD, weight initializations of the deep learning models, and data splits for datasets such as Coauthor Physics, Coauthor CS, BlogCatalog, and DBLP. For Cora, Citeseer, and Pubmed, we use predefined splits given in Yang et al. (2016).
@@
For the link prediction task, we rely on the results directly obtained from the paper for GVAE from Kipf and Welling (2016).
@@
Figure 2. Comparison of Learning Times (in seconds) i.e Across Datasets for HDGL and RelHD with GCN, GAT and SGC with Hyperparameter Tuning with Exploration of Search Space with 12 configurations. Runtime for HDGL and RelHD exclusively includes only time Elapsed for learning label representations using train/validation data.
[…]
5.5. Implementation Details
All experiments were conducted on Google Colab Pro with High-memory CPU. To ensure fair comparison, although not ideal for HDGL, we used the PyTorch framework (Paszke et al., 2019), complemented by the Deep Graph Library (DGL) (Wang et al., 2019) for importing datasets and implementing deep learning models. Due to the lack of support for bitwise majority operations in PyTorch, we convert our vectors to the bipolar space and utilize signed addition for compatibility. As mentioned in Section 3.2, HD computing works in both binary and bipolar spaces, allowing us to switch between them as necessary. None of our experiments used parallel execution or GPUs. We have also implemented PyTorch version of RelHD, as the experiments detailed in Kang et al. (2022) were reliant on hardware architecture implementations. We undertake this effort to compare with our model and use RelHD as a baseline in this study.